OpenStudy (anonymous):

Solve: x2 − 10x − 24 = 0. x = 2, x = 12 x = 2, x = −12 x = −2, x = −12 x = −2, x = 12

terenzreignz (terenzreignz):

Factor it

OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

or, since you have choices, just plug in the values for x, and see which ones actually work

OpenStudy (dls):

x2 − 10x − 24 = 0 x2-12x+2x-24=0 x(x-12)+2(x-12)=0 (x+2)(x-12)=0 x=-2 or x=12 Last option

OpenStudy (anonymous):

Oh . Okay thanks

OpenStudy (dls):

anytime =)

OpenStudy (compassionate):

how do you complete the square? x^2+10x+24=0 To solve this by completing the square method! x^2+10x+24=0 First thing is to apply any properties to cancel the constant term (24) on the left side of the equation. A property that came to mind and which is the best - is to subtract the constant (24) from both sides of the equation. We don't want the constant term on the left side of the equation as a process of completing the square method! (that is always the first step). Therefore by subtracting 24 from both sides of the equation.The equation becomes- x^2+10x+24-24= 0-24 x^2+10x = -24 The second step is to... find the constant term that makes the left side a perfect square trinomial by squaring half the coefficient of x. Add this constant to both sides of the equation.(This statement might be a little bit confusing but endeavor to sit down and make sense of it by reading it carefully and apply it)-The statement is no mistake at all. Okay lets apply step 2 rule- half of 10(co-efficient of x) is 5 and (5)^2 = 25.1.e (by squaring half the coefficient of x ) Add this 25 to both sides of the equation... x^2+10x = -24 x^2+10x+25 = -24+25 (x+5)(x+5) = 1 (x+5)^2 = 1 (x+5) = sqrt(1)) x+5 = +1 OR - 1 x+5= +1 x= 1-5 = -4

OpenStudy (dls):

lol

OpenStudy (hba):

the easy solution is x=-4 put it in the question